OpenStudy (anonymous):

bx-a=5a+c solve for x

jimthompson5910 (jim_thompson5910):

\[\Large bx-a=5a+c\] \[\Large bx=5a+c+a\] \[\Large bx=6a+c\] \[\Large x=\frac{6a+c}{b}\]

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

add "a" to both sides like this \[bx = 5a +c +a\] add like terms \[bx = 6a + c\] divide by "b" \[x = 6a + c \div b\]

jimthompson5910 (jim_thompson5910):

better to write \[x = (6a+c)\div b\] to make it clear that the entire quantity (6a+c) is divided by b

OpenStudy (anonymous):

Hey Jim how do you write the equation where it puts the denominator like the way you did.. ???

jimthompson5910 (jim_thompson5910):

type x=\frac{6a+c}{b}

jimthompson5910 (jim_thompson5910):

\frac{ allows for stuff like \[\Large \frac{x+y}{z}\]

OpenStudy (anonymous):

oh ok.. Thanks